Gummy Bears and Skittles are slightly differentiated products. The demand for a bag of Gummy Bears is qB = 80 – 5pB + 3pS, where pB is the price of a bag of Gummy Bears and pS is the price of a bag of Skittles. The demand for a bag of Skittles is qS = 50 – 4pS + 2pB. Under profitmaximizing Bertrand competition, find the equilibrium prices of both Gummy Bears and Skittles. Assume AC = MC = 10.
Answer #1

28 = 80-588 + 3P% TCB = 109. [sime AL=mc=107 TB = TRB - TCR In = P8 LB - 10qB PB-10798 (PB 10 ) (80-5P₂ + 3ps) TB = 80pg-5P₂

If you have any doubt feel free to ask.

Don't forget to thumbs up if it helped you.